is y=x^3-4x^2 symmetrical to origin
Answer by fractalier(6550)

To see if that is true, we compare f(-x) to -f(x)...if they are equal, then the function is symmetric about the origin...
Here, y = f(x) = x^3 - 4x^2.
f(-x) = -x^3 - 4x^2
-f(x) = -x^3 + 4x^2
Since they are not equal, the function is not symmetric about the origin.